Trying to help my son with his homework.
8(v-1)=24+3m
I really would like a breakdown to the answer please

Answer by stanbon(75887)

8(v-1)=24+3m
------
If you are asked to solve the equation for "v":
8(v-1) = 24 + 3m
----
multiply on the left side to get:
8v - 8 = 24 + 3m
----
Add 8 to both sides to get:
8v = 32 + 3m
----
Divide both sides by 8 to get:
v = 4 + (3/8)m
---------------------------------
If you are asked to solve for "m":
8(v-1) = 24 + 3m
----
multiply on the left side to get:
8v - 8 = 24 + 3m
----
Subtract 24 from both sides to get:
3m = 8v-32
---
Divide both sides by 3 to get:
m = (8/3)v - (32/3)
----------------------------
Cheers,
Stan H.
-----------------------------